# Discovering new robust local search algorithms with neuro-evolution

### v1

@@ 4 Computational experiments
In this section, we address four questions experimentally:
@@ 4.1 Experimental settings
Neuro-LS is implemented in Python 3.7 with Pytorch 1.4 library. [^3] The program source code, and benchmark instances are available at the url https://github.com/Salim-AMRI/NK_Landscape_Project.git. For all experiments with different values $N$ and $K$, we use the same architecture of the neural network composed of two hidden layers of size 10 and 5, with a total of $\mid\theta\mid=81$ parameters to calibrate when $d=1$ (when using observation functions $o_{1}$ and $o_{3}$) and $\mid\theta\mid=91$ parameters to calibrate when $d=2$ (when using observation functions $o_{2}$ and $o_{4}$).
@@ 4.2 Neuro-LS training phase
The evolution of the average score obtained on the Neuro-LS validation sets, considering as input the observation matrices $o_{1}$, $o_{2}$, $o_{3}$ and $o_{4}$ are respectively indicated with green, red, blue and yellow lines in Figure 3.
@@
The strategy, which obtains the best results on average over the validation set, is the strategy that uses the $o_{4}$ observation function. Remind that $o_{4}$ extracts information from both the rank of improving and deteriorating neighbors and the z-scores of fitness variations for each possible flip. This strategy learns faster than the other strategies and obtains more consistent scores across the different training runs.
@@ 4.3 Test phase
Table 1 summarizes the average score obtained by the four variants of Neuro-LS, with the different observation functions $o_{1}$, $o_{2}$, $o_{3}$ and $o_{4}$, and the three other competing methods, namely the ${\rm BHC}^{+}$, ${\rm FHC}^{+}$ and (1, $\lambda$)-ES algorithms, presented in section 3.2.
@@
Results underlined indicate significant better results obtained by a Neuro-LS strategy on average compared to all the baseline strategies (p-value below 0.001), measured with a Student t-test without assuming equal variance. [^5] The normality condition required for this test was first confirmed using a Shapiro statistical test on the empirical distributions of 100 iid scores obtained by each strategy.
@@
32 2 0.713 0.709 0.717 0.718 0.718 0.718 0.718
32 4 0.717 0.721 0.713 0.0.741 724 0.739 0.736
32 8 0.702 0.712 0.707 0.734 0.718 0.734 0.738
64 1 0.700 0.693 0.696 0.702 0.702 0.702 0.702
64 2 0.712 0.709 0.712 0.720 0.720 0.720 0.720
64 4 0.721 0.721 0.714 0.737 0.726 0.741 0.742
64 8 0.710 0.719 0.707 0.739 0.726 0.739 0.74
128 1 0.698 0.691 0.696 0.699 0.699 0.699 0.699
128 2 0.712 0.711 0.710 0.713 0.713 0.713 0.714
128 4 0.724 0.723 0.717 0.735 0.725 0.737 0.741
128 8 0.711 0.719 0.705 0.741 0.720 0.744 0.742
Table 1: Average results on test instances obtained by the baseline local search strategies and the Neuro-LS strategies with different type of observations given as input for different NK landscape configurations.
Table 1 shows that all the Neuro-LS variants always obtains better results than the baseline algorithms for all the configurations of NK landscape, but the difference in score is only really significant when $K=4$ and $K=8$ for all values of $N$. It means that our learned Neuro-LS strategies become more effective than other methods when the landscape is more rugged.
@@
Among the Neuro-LS strategies, we observe that the strategy using the second observation function $o_{2}$ gets the worst scores on the validation sets. This is because the neural network struggles to extract relevant information using the rawest possible information about the fitness of the current solution and the fitness of its neighbors.
The other strategies combined with the observation functions $o_{1}$, $o_{3}$, $o_{4}$ obtain almost the same score, when tested on instances from the same distribution as used in the training phase.
@@ 4.4 Out-of-distribution tests
We select the best Neuro-LS strategies trained on the distribution of NK landscape instances with size $N=64$ and $K=8$, and we test them on new instances of different sizes [^6] Note that a Neuro-LS strategy learn for a given size of problem can be applied for instances of different sizes, as the input of the neural-network $g_{\theta}$ does not depend on $N$ (see Section 3.1). of the Quadratic Unconstrained Binary Optimization (QUBO) problem.
The QUBO problem is a single objective pseudo-boolean optimization problem with quadratic interactions between binary variables. The objective function $f:\{0,1\}^{N}\to\mathbb{R}$ to maximize is defined by: $f(x)=x^{t}Qx$ where $Q$ is a real matrix of dimension $N\times N$ and $x^{t}$ is the transposed vector of $x$. We consider a black-box optimization scenario with an unknown matrix $Q$.
@@
Table 2 gives the experimental setup of PUBOi instances used in this work. We consider instances of size $N\in\{32,64,128,256\}$. The number of sub-functions $m$ tunes the density of the matrix ($16\%$ and $43\%$ for uniform instances respectively for the two values of $m$). Three types of interaction mechanisms are used between variables. The instances $I_{uni}$ have no specific important variables, i.e. $I_{uni}$ instances are similar to QUBO problems with a random matrix. The instances $I_{imp}$ have important variables: the marginal probability of having important variables is equal to $d=10$ times the probability of non-important ones. Additionally, for the $I_{ic}$ instances, the value of the co-appearance parameter is high, the selection of important variables is not independent, and the selection of important variables is concentrated. For each tuple of parameter values, we generate a test set $\mathcal{D}_{test}$ including $100$ independent instances.
@@
$(d,\alpha)$ (Degree of importance, $I_{uni}:(1,1)$, $I_{imp}:(10,1)$,
importance co-appearance) $I_{ic}:(10,1.09)$
@@
In Table 3, the best average result obtained among the above-mentioned methods for each configuration of QUBO test instances is highlighted in bold. Results underlined indicate significant better results obtained by a Neuro-LS strategy on average compared to all the baseline strategies (p-value below 0.001), measured with a Student t-test without assuming equal variance.
This table shows that Neuro-LS strategies that use observation functions $o_{3}$ and $o_{4}$, which are based on ranking neighboring solutions, get the best results. These strategies outperform not only the baseline algorithms but also the Neuro-LS strategies that rely on raw information about neighbors’ fitness (columns $o_{1}$ and $o_{2}$).
@@
When comparing results on columns $o_{3}$ and $o_{4}$, we observe that the introduction of additional z-score information in Neuro-LS produces better results for all instance types of larger size ($N=256$). As the choice of possible actions at each local search iteration becomes larger, using additional information on the fitness variations associated with each possible action seems beneficial. We will take a closer look at how the learned strategy exploits this information in the next section.
@@
$o_{1}$ $o_{2}$ $o_{3}$ $o_{4}$
32 $0.05$ $I_{uni}$ 60.8 59.3 59.4 1.3 33.8 61.1 61.0
32 $0.05$ $I_{imp}$ 43.8 43.4 44.1 3.7 32.4 44.6 43.1
32 $0.05$ $I_{ic}$ 41.6 41.5 41.0 -1.8 32.9 41.3 41.8
32 $0.20$ $I_{uni}$ 136.3 134.3 133.6 0.3 69.2 141.4 142.0
32 $0.20$ $I_{imp}$ 101.4 99.3 97.2 -1.4 67.7 103.3 100.7
32 $0.20$ $I_{ic}$ 92.8 92.0 91.2 3.8 68.7 93.5 91.9
64 $0.05$ $I_{uni}$ 188.0 186.7 187.1 4.5 64.9 193.0 192.6
64 $0.05$ $I_{imp}$ 138.8 135.8 136.5 -3.6 71.9 145.1 139.4
64 $0.05$ $I_{ic}$ 129.9 130.3 131.8 2.4 74.0 134.4 134.1
64 $0.20$ $I_{uni}$ 399.7 398.4 386.3 2.3 125.4 415.9 422.8
64 $0.20$ $I_{imp}$ 308.3 301.8 304.6 6.2 144.3 317.8 313.1
64 $0.20$ $I_{ic}$ 294.0 289.3 289.9 2.1 146.8 303.5 300.2
128 $0.05$ $I_{uni}$ 553.7 553.1 554.7 3.7 116.4 573.6 580.7
128 $0.05$ $I_{imp}$ 430.2 418.0 424.6 5.9 132.9 432.3 431.7
128 $0.05$ $I_{ic}$ 408.5 405.6 418.2 1.3 136.9 417.6 422.1
128 $0.20$ $I_{uni}$ 1167.7 1175.9 1136.2 -5.5 220.4 1212.1 1228.8
128 $0.20$ $I_{imp}$ 901.7 883.1 892.1 -7.9 269.6 928.1 937.6
128 $0.20$ $I_{ic}$ 893.3 889.7 878.8 -9.7 267.7 925.0 922.6
256 $0.05$ $I_{uni}$ 1636.5 1643.4 1628.3 4.9 188.8 1679.9 1699.6
256 $0.05$ $I_{imp}$ 1282.3 1244.3 1278.4 10.7 238.4 1287.5 1320.9
256 $0.05$ $I_{ic}$ 1228.3 1225.4 1244.6 13.3 239.9 1283.3 1292.3
256 $0.20$ $I_{uni}$ 3344.7 3367.0 3289.4 17.4 372.3 3474.5 3474.6
256 $0.20$ $I_{imp}$ 2584.4 2578.2 2577.2 16.2 433.8 2690.0 2709.0
256 $0.20$ $I_{ic}$ 2563.9 2617.5 2563.5 8.9 444.6 2697.2 2735.6
Table 3: Average score (fitness value) on test instances obtained by the baseline local search strategies and the Neuro-LS strategies for different QUBO problems of different sizes.
@@ 4.5.1 Emerging strategies for smooth landscape
Figure 4 displays a representative example of the trajectory performed by Neuro-LS with $N=64$ and $K=2$ and when using the observation vector $o_{1}(x)$ as input (vector of delta fitness).
@@
Figure 4: Fitness evolution curve and strategy used by Neuro-LS using $o_{1}$ observations for the resolution of an instance with smooth NK landscape ($N=64$ and $K=2$).
@@ 4.5.2 Emerging strategies for rugged landscape
For rugged landscapes, when $K=4$ and $K=8$, the emerging strategy is much more interesting. We detail here the most interpretable strategy learned using observation vectors $o_{1}$ and $o_{3}$, respectively corresponding to variations of fitness for each possible move and a ranking of the improving and deteriorating moves (see Section 2.2.2). The two other strategies using $o_{2}$ and $o_{4}$ observation functions are described in Appendix A.
@@
Figure 5 displays a representative example of the trajectory performed by Neuro-LS using observation $o_{1}$ when $N=64$ and $K=8$. On this plot, we observe that the emerging Neuro-LS strategy has two successive operating modes:
@@
Figure 5: Fitness evolution curve and strategy used by Neuro-LS using $o_{3}$ observations for the resolution of an instance with significantly rugged NK landscape ($N=64$ and $K=8$).
To explain why Neuro-LS has this behaviour, we display in Figure 6 the continuous function $g_{\theta}$ learned by the neural network. It explains why Neuro-LS chooses one movement over another. Indeed, as seen in Section 3.1, given the vector of observation, which in this case is equal to $o_{1}(x)=(\Delta_{1}(x),\dots,\Delta_{N}(x))$, the neural network outputs a vector $g_{\theta}(o)=(g_{\theta}(\Delta_{1}(x)),\dots,g_{\theta}(\Delta_{N}(x))\in\mathbb{R}^{N}$. Then, the action $a=\operatorname*{argmax}_{i\in\llbracket 1,N\rrbracket}g_{\theta}(o_{i})$ is chosen by Neuro-LS. Thus, the output of the neural network, $g_{\theta}(\Delta_{i}(x))$, can be interpreted as a preference score for a move associated to a variation of fitness $\Delta_{i}(x)$.
@@
As seen in Figure 6, the preference score is greater for moves that corresponds to a small variation of fitness (in green), which explains the first operating mode mentioned earlier, with Neuro-LS performing only small steps and avoiding large improvement of fitness, which may cause to get trap too early in a local maximum. When there are only movements with negative score improvements, we see that $g_{\theta}$ gives a higher preference score for the most deteriorating move, which explains the second operating mode, when Neuro-LS performs a jump, when there are no more score-improving moves.
Note that this strategy depends on the amplitude of the variations of fitness of the problem. If the fitness function undergoes a change of scale, the strategy learned by Neuro-LS is completely disrupted. This explains why the same Neuro-LS strategy using the $o_{1}$ observation vector does not score well when applied to the UBQP pseudo-Boolean problem, which does not have the same fitness variation amplitudes (see Table 3). We will see in the following that using an information based on the ranked allows to avoid this problem.
@@
Figure 7 displays a representative example of the trajectory performed by Neuro-LS using observation $o_{3}$ when $N=64$ and $K=8$.
Figure 7: Fitness evolution curve and strategy used by Neuro-LS using $o_{3}$ observations for the resolution of an instance with significantly rugged NK landscape ($N=64$ and $K=8$).
@@
To explain why Neuro-LS has this behaviour, we display in Figure 8 the continuous function $g_{\theta}$ learned by the neural network. We remain the reader that the values $o_{3}(x)_{i}$ given as input to the neural network corresponds to a ranking of the improvement moves, which receive a strictly positive value in the range ]0,1]]0,1]] 0 , 1 ], and the deteriorating moves, which receive a strictly negative value in the range [1,0[[-1,0[[ - 1 , 0 [. We see on this figure that Neuro-LS gives the higher preference score for a median improving move (with a value around 0.5), which explain the median hill climbing behavior observed during the trajectory. When there is only negative moves, Neuro-LS prefers the worst deteriorating move with a value of -1, which explain the jump when there are no more improving moves.
@@
Figure 8: Output of the neural network $g_{\theta}(o_{3}(x)_{i}))$ (y-axis) for each input $o_{3}(x)_{i}$ (x-axis). It corresponds to the values collected during 10 trajectories of Neuro-LS on the test instances of the NK Landscape problem with $N=64$ and $K=8$. ).

### v2

@@ 4 Computational experiments
In this section, we target four aspects of Neuro-LS experimentally:
@@ 4.1 Experimental settings
Neuro-LS is implemented in Python 3.7 with Pytorch 1.4 library. [^5] The program source code, and benchmark instances are available at the url https://github.com/Salim-AMRI/NK_Landscape_Project.git. For all experiments with different values $N$ and $K$, we use the same architecture of the neural network composed of two hidden layers of size 10 and 5, with a total of $\mid\theta\mid=81$ parameters to calibrate when $d=1$ (when using observation functions $o^{1}$ and $o^{3}$) and $\mid\theta\mid=91$ parameters to calibrate when $d=2$ (when using observation functions $o^{2}$ and $o^{4}$).
@@ 4.2 Neuro-LS training phase
The evolution of the average score obtained on the Neuro-LS validation sets, considering as input the observation matrices $o^{1}$, $o^{2}$, $o^{3}$ and $o^{4}$ are respectively indicated with green, red, blue and yellow lines in Figure 3.
@@
The strategy, which obtains the best results on average over the validation set, is the strategy that uses the $o^{4}$ observation function (let us recall that $o^{4}$ extracts information from both the rank of improving and deteriorating neighbors and the z-scores of fitness variations for each possible flip). This strategy learns faster than the other strategies and obtains more consistent scores across the different training runs.
@@ 4.3 Test phase
Table 1 summarizes the average score obtained by the four variants of Neuro-LS, with the different observation functions $o^{1}$, $o^{2}$, $o^{3}$ and $o^{4}$, and the three other competing methods, namely the ${\rm BHC}^{+}$, ${\rm FHC}^{+}$ and (1, $\lambda$)-ES algorithms, presented in Section 3.2.
@@
Results underlined indicate significant better results obtained by a Neuro-LS strategy on average compared to all the baseline strategies (p-value below 0.001), measured with a one-tailed Student t-test without assuming equal variance and with a Bonferroni correction to correct the significance threshold for multiple comparisons. [^7] The normality condition required for this test was first confirmed using a Shapiro statistical test on the empirical distributions of 100 iid scores obtained by each strategy.
@@
32 2 0.716 0.714 0.717 0.719 0.710 0.719 0.719
32 4 0.717 0.723 0.713 0.742 0.724 0.739 0.736
32 8 0.705 0.716 0.707 0.730 0.718 0.731 0.729
64 1 0.700 0.694 0.696 0.701 0.699 0.701 0.701
64 2 0.716 0.713 0.712 0.717 0.717 0.717 0.717
64 4 0.718 0.726 0.714 0.743 0.726 0.742 0.744
64 8 0.706 0.714 0.707 0.737 0.726 0.739 0.738
128 1 0.699 0.694 0.696 0.699 0.699 0.699 0.699
128 2 0.715 0.709 0.710 0.716 0.713 0.716 0.716
128 4 0.723 0.726 0.717 0.738 0.726 0.740 0.740
128 8 0.710 0.721 0.705 0.739 0.720 0.740 0.743
Table 1: Average results on test instances obtained by the baseline local search strategies and the Neuro-LS strategies with different type of observations given as input for different NK landscape configurations. Best result on each line is written in bold. Results underlined indicate significant better results obtained by a Neuro-LS strategy on average compared to all the baseline strategies (p-value below 0.001), measured with a one-tailed Student t-test without assuming equal variance and with a Bonferroni correction to correct the significance threshold for multiple comparisons.
Table 1 shows that all the Neuro-LS variants always obtains equal or better results than the baseline algorithms for all the configurations of NK landscape, but the difference in score is only really significant when $K=4$ and $K=8$ for all values of $N$. It means that our learned Neuro-LS strategies become more effective than other methods when the landscape is more rugged.
@@
Among the Neuro-LS strategies, we observe that the strategy using the second observation function $o^{2}$ gets the worst scores on the validation sets. This is because the neural network struggles to extract relevant information using the rawest possible information about the fitness of the current solution and the fitness of its neighbors.
The other strategies combined with the observation functions $o^{1}$, $o^{3}$, $o^{4}$ obtain almost the same score, when tested on instances from the same distribution as used in the training phase.
@@ 4.4 Out-of-distribution tests
We select the best Neuro-LS strategies trained on the distribution of NK landscape instances with size $N=64$ and $K=8$, and we test them on new instances of different sizes [^8] Note that a Neuro-LS strategy learned for a given size of problem can be applied for instances of different sizes, as the input of the neural-network $g_{\theta}$ does not depend on $N$ (see Section 3.1). of the Quadratic Unconstrained Binary Optimization (QUBO) problem.
The QUBO problem is a single objective pseudo-Boolean optimization problem with quadratic interactions between binary variables. The objective function $f:\{0,1\}^{N}\to\mathbb{R}$ to maximize is defined by: $f(x)=x^{\intercal}Qx$ where $Q$ is a real matrix of dimension $N\times N$ and $x^{\intercal}$ is the transposed vector of $x$. We consider a black-box optimization scenario with an unknown matrix $Q$.
@@
Table 2 gives the experimental setup of PUBOi instances used in this work. We consider instances of size $N\in\{32,64,128,256\}$. The number of sub-functions $m$ tunes the density of the matrix ($16\%$ and $43\%$ for uniform instances respectively for the two values of $m$). Three types of interaction mechanisms are used between variables. The instances $I_{uni}$ have no specific important variables, i.e. $I_{uni}$ instances are similar to QUBO problems with a random matrix. The instances $I_{imp}$ have important variables: the marginal probability of having important variables is equal to $v=10$ times the probability of non-important ones. Additionally, for the $I_{ic}$ instances, the value of the co-appearance parameter is high, the selection of important variables is not independent, and the selection of important variables is concentrated. For each tuple of parameter values, we generate a test set $\mathcal{D}_{test}$ including $100$ independent instances.
@@
$(v,\alpha)$ (Degree of importance, $I_{uni}:(1,1)$, $I_{imp}:(10,1)$,
co-appearance parameter) $I_{ic}:(10,1.09)$
@@
In Table 3, the best average result obtained among the above-mentioned methods for each configuration of QUBO test instances is highlighted in bold. Results underlined indicate significant better results obtained by a Neuro-LS strategy on average compared to all the baseline strategies (p-value below 0.001), measured with a one-tailed Student t-test without assuming equal variance and with a Bonferroni correction to correct the significance threshold for multiple comparisons.
This table shows that Neuro-LS strategies that use observation functions $o^{3}$ and $o^{4}$, which are based on ranking neighboring solutions, get the best results. These strategies outperform not only the baseline algorithms but also the Neuro-LS strategies that rely on raw information about neighbors’ fitness (columns $o^{1}$ and $o^{2}$).
@@
When comparing results on columns $o^{3}$ and $o^{4}$, we observe that the introduction of additional z-score information in Neuro-LS produces better results for all instance types of larger size ($N=256$). As the choice of possible actions at each local search iteration becomes larger, using additional information on the fitness variations associated with each possible action seems beneficial. We will take a closer look at how the learned strategy exploits this information in the next section.
@@
$o^{1}$ $o^{2}$ $o^{3}$ $o^{4}$
32 $0.05$ $I_{uni}$ 58.7 59.0 58.7 1.1 12.4 60.6 61.7
32 $0.05$ $I_{imp}$ 44.4 42.4 43.2 2.7 10.9 44.7 43.6
32 $0.05$ $I_{ic}$ 41.7 41.5 41.3 -0.18 5.3 41.1 41.7
32 $0.20$ $I_{uni}$ 131.4 133.5 131.4 -0.75 85.1 139.6 144.3
32 $0.20$ $I_{imp}$ 101.8 98.4 99.8 -4.9 73.2 103.8 100.7
32 $0.20$ $I_{ic}$ 92.9 93.2 92.9 -2.9 74.1 93.7 92.0
64 $0.05$ $I_{uni}$ 187.1 184.6 183.7 -2.87 40.8 195.4 194.4
64 $0.05$ $I_{imp}$ 139.8 135.0 136.5 -3.3 31.6 141.1 141.0
64 $0.05$ $I_{ic}$ 131.1 129.4 130.4 -0.85 25.9 134.4 132.5
64 $0.20$ $I_{uni}$ 396.8 391.3 393.8 6.36 133.9 418.2 415.9
64 $0.20$ $I_{imp}$ 310.7 305.1 312.6 -2.78 201.4 319.9 312.7
64 $0.20$ $I_{ic}$ 295.5 292.4 293.0 2.46 198.3 303.9 301.1
128 $0.05$ $I_{uni}$ 546.0 556.9 552.1 5.5 144.7 572.8 583.3
128 $0.05$ $I_{imp}$ 423.1 414.0 418.9 -8.3 75.3 432.5 432.4
128 $0.05$ $I_{ic}$ 406.0 402.0 405.1 3.2 132.1 420.4 424.1
128 $0.20$ $I_{uni}$ 1156.7 1171.4 1123.0 -13.4 338.1 1215.1 1228.0
128 $0.20$ $I_{imp}$ 901.6 899.1 895.6 -8.97 256.5 923.3 919.9
128 $0.20$ $I_{ic}$ 888.3 893.7 881.1 11.7 399.4 917.1 914.0
256 $0.05$ $I_{uni}$ 1620.7 1630.6 1609.5 7.8 438.8 1681.3 1701.0
256 $0.05$ $I_{imp}$ 1275.0 1252.6 1275.8 -4.12 342.4 1299.5 1320.9
256 $0.05$ $I_{ic}$ 1226.6 1230.6 1235.3 1.38 262.1 1276.6 1295.1
256 $0.20$ $I_{uni}$ $3$.6 3376.1 3282.9 -13.3 1083.8 3485.0 3484.6
256 $0.20$ $I_{imp}$ 2604.0 2580.3 2574.3 -20.4 693.9 2687.0 2697.6
256 $0.20$ $I_{ic}$ 2592.9 2574.4 2584.4 8.3 868.5 2698.6 2713.4
Table 3: Average score (fitness value) on test instances obtained by the baseline local search strategies and the Neuro-LS strategies for different QUBO problems of different sizes. Best result on each line is written in bold. Results underlined indicate significant better results obtained by a Neuro-LS strategy on average compared to all the baseline strategies (p-value below 0.001), measured with a one-tailed Student t-test without assuming equal variance and with a Bonferroni correction to correct the significance threshold for multiple comparisons.
@@ 4.5.1 Emerging strategies for smooth landscape
Figure 4 displays a representative example of the trajectory performed by Neuro-LS with $N=64$ and $K=2$ and when using the observation vector $o^{1}(x)$ as input (vector of delta fitness).
@@
Figure 4: Fitness evolution curve and strategy used by Neuro-LS using $o^{1}$ observations for the resolution of an instance with smooth NK landscape ($N=64$ and $K=2$).
@@ 4.5.2 Emerging strategies for rugged landscape
For rugged landscapes, when $K=4$ and $K=8$, the emerging strategy is much more interesting. We detail here the most interpretable strategy learned using observation vectors $o^{1}$ and $o^{3}$, respectively corresponding to variations of fitness for each possible move and a ranking of the improving and deteriorating moves (see Section 2.2.2). The two other strategies using $o^{2}$ and $o^{4}$ observation functions are described in Appendix A.
@@
Figure 5 displays a representative example of the trajectory performed by Neuro-LS using observation $o^{1}$ when $N=64$ and $K=8$. On this plot, we observe that the emerging Neuro-LS strategy has two successive operating modes:
@@
Figure 5: Fitness evolution curve and strategy used by Neuro-LS using $o^{3}$ observations for the resolution of an instance with significantly rugged NK landscape ($N=64$ and $K=8$).
To explain why Neuro-LS has this behaviour, we display in Figure 6 the continuous function $g_{\theta}$ learned by the neural network. It explains why Neuro-LS chooses one movement over another. Indeed, as seen in Section 3.1, given the vector of observation, which in this case is equal to $o^{1}(x)=(\Delta_{1}(x),\dots,\Delta_{N}(x))$, the neural network outputs a vector $g_{\theta}(o)=(g_{\theta}(\Delta_{1}(x)),\dots,g_{\theta}(\Delta_{N}(x))\in\mathbb{R}^{N}$. Then, the action $a=\operatorname*{argmax}_{i\in\llbracket 1,N\rrbracket}g_{\theta}(o_{i})$ is chosen by Neuro-LS. Thus, the output of the neural network, $g_{\theta}(\Delta_{i}(x))$, can be interpreted as a preference score for a move associated to a variation of fitness $\Delta_{i}(x)$.
@@
As seen in Figure 6, the preference score is greater for moves that corresponds to a small variation of fitness (in green), which explains the first operating mode mentioned earlier, with Neuro-LS performing only small steps and avoiding large improvement of fitness, which may cause it to get trapped too early in a local maximum. When there are only movements with negative score improvements, we see that $g_{\theta}$ gives a higher preference score for the most deteriorating move, which explains the second operating mode, when Neuro-LS performs a jump, when there are no more score-improving moves.
Note that this strategy depends on the amplitude of the variations of fitness of the problem. If the fitness function undergoes a change of scale, the strategy learned by Neuro-LS is completely disrupted. This explains why the same Neuro-LS strategy using the $o^{1}$ observation vector does not score well when applied to the UBQP pseudo-Boolean problem, which does not have the same fitness variation amplitudes (see Table 3). We will see in the following that using an information based on the ranked allows to avoid this problem.
@@
Figure 7 displays a representative example of the trajectory performed by Neuro-LS using observation $o^{3}$ when $N=64$ and $K=8$.
Figure 7: Fitness evolution curve and strategy used by Neuro-LS using $o^{3}$ observations for the resolution of an instance with significantly rugged NK landscape ($N=64$ and $K=8$).
@@
To explain why Neuro-LS has this behaviour, we display in Figure 8 the continuous function $g_{\theta}$ learned by the neural network. We remain the reader that the values $o^{3}(x)_{i}$ given as input to the neural network corresponds to a ranking of the improvement moves, which receive a strictly positive value in the range ]0,1]]0,1]] 0 , 1 ], and the deteriorating moves, which receive a strictly negative value in the range [1,0[[-1,0[[ - 1 , 0 [. We see on this figure that Neuro-LS gives the higher preference score for a median improving move (with a value around 0.5), which explain the median hill climbing behavior observed during the trajectory. When there is only negative moves, Neuro-LS prefers the worst deteriorating move with a value of -1, which explain the jump when there are no more improving moves.
@@
Figure 8: Output of the neural network $g_{\theta}(o^{3}(x)_{i}))$ (y-axis) for each input $o^{3}(x)_{i}$ (x-axis). It corresponds to the values collected during 10 trajectories of Neuro-LS on the test instances of the NK Landscape problem with $N=64$ and $K=8$. ).